OpenStudy (anonymous):

What is the solution to 3|x-1| >= 12?

OpenStudy (texaschic101):

3 | x - 1 | > = 12 -- divide both sides by 3 | x - 1 | >= 4 x - 1 >= 4 x >= 4 + 1 x >= 5 x - 1 <= -4 x <= -4 + 1 x <= -3

OpenStudy (texaschic101):

read and learn... http://www.purplemath.com/modules/absineq.htm

OpenStudy (anonymous):

thank you!

OpenStudy (texaschic101):

sure thing :)